A 250-turn rectangular loop that is 26 cm long and 17 cm wide is placed in a 0.50-T magnetic field. If the maximum torque is 21 N · m, determine the current in the loop.
Answer #1

No of turns, N = 250.

Area of a single turn, A = 26cm X 17cm = 442 cm2 = 0.0442 m2

Magnetic flux density , B = 0.50 T

Current in the loop = I ( assume)

Maximum torque. T max = 21 N.m

We know that Tmax = N I BA

I = (Tmax / NBA) = ( 21 / ( 250 * 0.0442* 0.50)) = 3.80 A (ans)